# Coloring and density theorems for configurations of a given volume

### v2

@@ 9. Large rectangular boxes: proof of Theorem 5
Fix $n\geqslant m+1$. If the statement in part (a) of the theorem did not hold, then there would exist a sequence of positive numbers $(\lambda_{j})_{j=1}^{\infty}$ converging to $\infty$ such that for every index $j$ the set $A$ does not contain all vertices of any $m$-dimensional rectangular box (aligned as claimed in the theorem) in $\mathbb{R}^{n}$ with volume precisely $\lambda_{j}^{m}$. Passing to a subsequence we could also ensure $\lambda_{j+1}\geqslant 2\lambda_{j}$ for each index $j$. The usual localization argument will allow us to work on a sufficiently large cube $[0,R]^{n}$ and with sufficiently (but finitely) many scales $\lambda_{j}$, $1\leqslant j\leqslant J$. Indeed, leaving the choice of $J$ (depending on $\bar{\delta}_{n}(A)$) for later, we can find a number $R$ larger than $\lambda_{J}$ for which there exists $x\in\mathbb{R}^{n}$ such that
@@
The set $A^{\prime}:=(A-x)\cap[0,R]^{n}$ is then a subset of the cube $[0,R]^{n}$ with density at least $\delta$ and it still does not contain the required boxes of volumes $\lambda_{j}^{m}$, $1\leqslant j\leqslant J$.
Thus, it is sufficient to work, from the very beginning, with a measurable subset $A$ of a large cube $[0,R]^{n}$ with density $\delta>0$. We will obtain a contadiction with the assumption that the $m$-volumes of the (appropriately aligned) boxes with vertices in $A$ omit the numbers $\lambda_{j}^{m}$, $1\leqslant j\leqslant J$. Here we work with a large number $J$ that we can choose depending on $\delta$, but the cube side-length $R>0$ needs to be an arbitrary number larger than some threshold (which in turn is allowed to depend on $\delta$ and $J$).
@@
The analogue of Lemma 8.1 holds and it is shown in exactly the same way. We will apply the decomposition (8.3) here as well. The following three lemmas respectively handle the structured, uniform, and error parts.
@@
The lower bound that we prove is in fact a concrete one, involving a certain power of $\delta$, but Theorem 5 is not quantitative anyway.
@@
$E$ :=$:=\big{\{}(u,v,x,y)\in D:|u_{k}|<\theta\lambda\text{ for some }1\leqslant k\leqslant m-1\big{\}},$
@@
Split $[0,\lceil R/\lambda\rceil\lambda]\supseteq[0,R]$ into intervals of length $\lambda$, denoting the resulting partition by $\mathcal{I}$, and split $[0,\lceil R/\lambda\rceil\lambda]^{n-m+1}\supseteq[0,R]^{n-m+1}$ into cubes of side-length $\lambda$, denoting the obtained partition by $\mathcal{Q}$. Then
@@
×$\times R^{-n}\lambda^{-m+1}(\sigma\ast\mathbbm{k}_{t})_{a}(v)\Big{(}\prod_{k=1}^{m-1}\mathbbm{1}_{[-\lambda,-\theta\lambda]\cup[\theta\lambda,\lambda]}(u_{k})\Big{)}\,\textup{d}v\,\textup{d}y\,\textup{d}u\,\textup{d}x\,\frac{\textup{d}t}{2\pi t}.$
@@
$\lesssim\int_{\tau}^{\varepsilon}t^{-1/2}\,\textup{d}t\lesssim\varepsilon^{1/2}$
@@
×$\qquad\times R^{-n}\lambda^{-m+1}(\sigma\ast\mathbbm{k}_{t})_{a}(v)\Big{(}\prod_{k=1}^{m-1}\mathbbm{1}_{[-\lambda,-\theta\lambda]\cup[\theta\lambda,\lambda]}(u_{k})\Big{)}\,\textup{d}v\,\textup{d}y\,\textup{d}u\,\textup{d}x\,\frac{\textup{d}s}{s}\,\frac{\textup{d}t}{2\pi t},$
@@
$-\int_{\mathbb{R}^{n-m+1}}\int_{\mathbb{R}^{n-m+1}}f(x,y;u)f(x,y+v;u)(\sigma\ast\mathbbm{k}_{t})_{a}(v)\,\textup{d}v\,\textup{d}y$
@@
×$\times(\sigma_{a}\ast\mathbbm{g}_{r})(z-z^{\prime})\,\lambda^{-m+1}\Big{(}\prod_{k=1}^{m-1}\mathbbm{1}_{[-\lambda,-\theta\lambda]\cup[\theta\lambda,\lambda]}(u_{k})\Big{)}\,\textup{d}z\,\textup{d}z^{\prime}\,\textup{d}u\,\textup{d}x\,\frac{\textup{d}s}{s}\,\frac{\textup{d}t}{t}.$
@@
Due to our assumption on the exponential growth of the numbers $\lambda_{j}$, every positive number $s$ lies in a bounded number of intervals $[\varepsilon\lambda_{j}/(10e),\lambda_{j}/10]$, $1\leqslant j\leqslant J$, with a bound depending on $\varepsilon$ alone. Therefore, applying the last estimate with $\lambda=\lambda_{j}$ and summing in $j$, we get
@@
×$\times\Big{(}\prod_{\begin{subarray}{c}1\leqslant i\leqslant m\\ i\neq k\end{subarray}}\mathbbm{g}_{s\gamma_{i}}(x_{i}^{0}-x_{i}^{1})\Big{)}\,\textup{d}(x_{1}^{0},x_{1}^{1},x_{2}^{0},x_{2}^{1},\ldots,x_{m}^{0},x_{m}^{1})\,\frac{\textup{d}s}{s}$
for any $1\leqslant k\leqslant m$, any parameters $\gamma_{1},\ldots,\gamma_{m}\in(0,\infty)$, and a real-valued bounded and compactly supported measurable function $F$ on $\mathbb{R}^{d_{1}+d_{2}+\cdots+d_{m}}$. In our case, $d_{1}=\cdots=d_{m-1}=1$ and $d_{m}=n-m+1$, but the greater generally will only make the arguments more elegant and symmetric.
@@
::=:= $\prod_{(r_{1},\ldots,r_{k-1},r_{k+1},\ldots,r_{n})\in\{0,1\}^{m-1}}F(x_{1}^{r_{1}},\ldots,x_{k-1}^{r_{k-1}},x,x_{k+1}^{r_{k+1}},\ldots,x_{m}^{r_{m}})$
@@
$00=\mathcal{N}_{\lambda_{j}}^{0}(A;R)\geqslant\frac{1}{3}\mathcal{N}_{\lambda_{j}}^{1}(A;R)>$ 0
@@ 10. Hyperbolic embeddings: proof of Theorem 10
Rotating the plane by $45^{\circ}$ and stretching it by the factor $\sqrt{2}$ we can, more conveniently, replace the hyperbola $\mathscr{H}$ with the rotated hyperbola
@@
Here $u$ is a shorter notation for the $n$-tuple $(u_{1},\ldots,u_{n})$. As expected, $\mathcal{N}_{\lambda}^{0}(A;R)$ can again be obtained in the limit of $\mathcal{N}_{\lambda}^{\varepsilon}(A;R)$ as $\varepsilon\to 0$, which is shown easily, as in Lemma 8.1. For every $a,\lambda>0$ the support of the anisotropically dilated measure $\mu_{a\lambda,a\lambda^{-1}}$ lies on the hyperbola $\mathscr{H}^{\prime}$ stretched by the factor $a$. Thus, $\mathcal{N}_{\lambda}^{0}(A;R)>0$ for any value of $\lambda>0$ clearly means that there exists the desired configuration (1.5) inside the set $A$.
@@
$[0,\lceil R/\lambda\rceil\lambda]\times[0,\lceil R\lambda\rceil\lambda^{-1}]\supseteq[0,R]^{2}$
@@
×$\times(\mu\ast\mathbbm{k}_{t})_{a_{j}\lambda,a_{j}\lambda^{-1}}(u_{j})\Big{(}\prod_{\begin{subarray}{c}1\leqslant k\leqslant n\\ k\neq j\end{subarray}}(\mu\ast\mathbbm{g}_{t})_{a_{k}\lambda,a_{k}\lambda^{-1}}(u_{k})\Big{)}\,\textup{d}u\,\textup{d}z\,\frac{\textup{d}t}{2\pi t}.$
@@
×$\times(\mu\ast\mathbbm{k}_{t})_{a_{j}\lambda,a_{j}\lambda^{-1}}(u_{j})\Big{(}\prod_{\begin{subarray}{c}1\leqslant k\leqslant n\\ k\neq j\end{subarray}}(\mu\ast\mathbbm{g}_{t})_{a_{k}\lambda,a_{k}\lambda^{-1}}(u_{k})\Big{)}\,\textup{d}u\,\textup{d}z\,\frac{\textup{d}t}{2\pi t}.$
@@
×$\qquad\qquad\quad\times(\mu\ast\mathbbm{g}_{t/\sqrt{2}})_{a_{j}\lambda,a_{j}\lambda^{-1}}(w-w^{\prime})\,\textup{d}w\,\textup{d}w^{\prime},$
@@
×$\times(\mu\ast\mathbbm{g}_{t/\sqrt{2}})_{a_{j}\lambda,a_{j}\lambda^{-1}}(w-w^{\prime})\Big{(}\prod_{\begin{subarray}{c}1\leqslant k\leqslant n\\ k\neq j\end{subarray}}(\mu\ast\mathbbm{g}_{t})_{a_{k}\lambda,a_{k}\lambda^{-1}}(u_{k})\Big{)}\,\textup{d}w\,\textup{d}w^{\prime}\,\textup{d}(u_{k})_{k\neq j}\,\frac{\textup{d}t}{t}$
@@
×$\times(\mu\ast\mathbbm{g}_{t/\sqrt{2}})_{a_{j}\lambda,a_{j}\lambda^{-1}}(w-w^{\prime})\Big{(}\prod_{\begin{subarray}{c}1\leqslant k\leqslant n\\ k\neq j\end{subarray}}(\mu\ast\mathbbm{g}_{t})_{a_{k}\lambda,a_{k}\lambda^{-1}}(u_{k})\Big{)}\,\textup{d}w\,\textup{d}w^{\prime}\,\textup{d}(u_{k})_{k\neq j}\,\frac{\textup{d}t}{t}.$
@@ 11. Remarks for harmonic analysts
One can wonder about an analytical reason why a density result for area $1$ rectangles is not possible in any dimension, while we were able to prove such a result for area 1 right-angled triangles.
Namely, take Schwartz functions $\varphi$ and $\psi$, respectively defined on $\mathbb{R}^{d_{1}}$ and $\mathbb{R}^{d_{2}}$, that have compact frequency supports and suppose that at least one of these supports does not contain the origin. Also, take some $4$-tuple $(p_{1},p_{2},p_{3},p_{4})$ of Lebesgue exponents in $[1,\infty]$ satisfying the Hölder scaling,
@@
×$\times\varphi_{t}(x-x^{\prime})\psi_{1/t}(y-y^{\prime})\,\textup{d}y\,\textup{d}y^{\prime}\,\textup{d}x\,\textup{d}x^{\prime}\,\frac{\textup{d}t}{t}\Big{|}\leqslant C_{r,R}\prod_{k=1}^{4}\|f_{k}\|_{\textup{L}^{p_{k}}(\mathbb{R}^{d_{1}+d_{2}})}$
grows like a multiple of $\log(R/r)$ as $R/r\to\infty$. This is not any better than in the trivial estimate obtained from Hölder’s inequality applied for each fixed $t$. Instead of giving a direct argument for this observation, one could argue that any nontrivial cancellation, i.e., a bound of the form $o(\log(R/r))$ as $R/r\to\infty$, would translate to a positive result for rectangles of area $1$. We would simply adjust the proof of Theorem 5 to work for fixed area rectangles, making it similar to the proof of Theorem 1, which we know to be impossible due to Theorems 3 and 4.
@@
so we only need to make sure that $\varphi$ and $\psi$ are such that the last integral is nonzero. The fact that the above quadrilinear form does not satisfy any nontrivial estimates is in a sharp contrast with results for the so-called twisted paraproduct, where the scales are aligned as $\varphi_{t},\psi_{t}$, rather than $\varphi_{t},\psi_{1/t}$. It was shown to be bounded, at least when the exponents $p_{k}$ are greater than $2$, by Durcik [18, 19] and, in a dyadic model, previously also by the author [39].
@@
now actually has non-trivial cancellation over the scales $t$. Its best constants $C^{\prime}_{r,R}$ grow at most like $O((\log(R/r))^{1/2})$, which is easily seen similarly as in the proof of Lemma 8.4: by separating one of the functions and applying the classical square function bound to it. However, one could still ask if the above trilinear form is also bounded as $r\to 0+$ and $R\to\infty$, or, equivalently, if the constants $C^{\prime}_{r,R}$ are even independent of the scale truncation parameters $r$ and $R$. If, on the other hand, the scales were aligned as $\varphi_{t},\psi_{t}$, this would again be the bound for the twisted paraproduct and it would follow from [39, Theorem 1].

### v3

@@ 9. Large rectangular boxes: proof of Theorem 5
Fix $n\geqslant m+1$. If the statement in part (a) of the theorem did not hold, then there would exist a sequence of positive numbers $(\lambda_{j})_{j=1}^{\infty}$ converging to $\infty$ such that for every index $j$ the set $A$ does not contain all vertices of any $m$-dimensional rectangular box (aligned as claimed in the theorem) in $\mathbb{R}^{n}$ with volume precisely $\lambda_{j}^{m}$. Passing to a subsequence we could also ensure $\lambda_{j+1}\geqslant 2\lambda_{j}$ for each index $j$. The usual localization argument will allow us to work on a sufficiently large cube $0,R]^{n}$ and with sufficiently (but finitely) many scales $\lambda_{j}$, $1\leqslant j\leqslant J$. Indeed, leaving the choice of $J$ (depending on $\bar{\delta}_{n}(A)$) for later, we can find a number $R$ larger than $\lambda_{J}$ for which there exists $x\in\mathbb{R}^{n}$ such that
@@
The set $A^{\prime}:=(A-x)\cap[0,R]^{n}$ is then a subset of the cube $0,R]^{n}$ with density at least $\delta$ and it still does not contain the required boxes of volumes $\lambda_{j}^{m}$, $1\leqslant j\leqslant J$.
Thus, it is sufficient to work, from the very beginning, with a measurable subset $A$ of a large cube $0,R]^{n}$ with density $\delta>0$. We will obtain a contradiction with the assumption that the $m$-volumes of the (appropriately aligned) boxes with vertices in $A$ omit the numbers $\lambda_{j}^{m}$, $1\leqslant j\leqslant J$. Here we work with a large number $J$ that we can choose depending on $\delta$, but the cube side-length $R>0$ needs to be an arbitrary number larger than some threshold (which in turn is allowed to depend on $\delta$ and $J$).
@@
The analogue of Lemma 8.1 holds and is shown in exactly the same way. We will apply the decomposition (8.3) here as well. The following three lemmas respectively handle the structured, uniform, and error parts.
@@
The lower bound that we prove is, in fact, a concrete one, involving a certain power of $\delta$, but Theorem 5 is not quantitative anyway.
@@
Split $0,\lceil R/\lambda\rceil\lambda]\supseteq[0,R]$ into intervals of length $\lambda$, denoting the resulting partition by $\mathcal{I}$, and split $0,\lceil R/\lambda\rceil\lambda]^{n-m+1}\supseteq[0,R]^{n-m+1}$ into cubes of side-length $\lambda$, denoting the obtained partition by $\mathcal{Q}$. Then
@@
$\times R^{-n}\lambda^{-m+1}(\sigma\ast\mathbbm{k}_{t})_{a}(v)\Big(\prod_{k=1}^{m-1}\mathbbm{1}_{[-\lambda,-\theta\lambda]\cup[\theta\lambda,\lambda]}(u_{k})\Big)\,\textup{d}v\,\textup{d}y\,\textup{d}u\,\textup{d}x\,\frac{\textup{d}t}{2\pi t}.$
@@
$\lesssim\int_{\tau}^{\varepsilon}t^{-1/2}\,\textup{d}t\lesssim\varepsilon^{1/2}$ .
@@
$\qquad\times R^{-n}\lambda^{-m+1}(\sigma\ast\mathbbm{k}_{t})_{a}(v)\Big(\prod_{k=1}^{m-1}\mathbbm{1}_{[-\lambda,-\theta\lambda]\cup[\theta\lambda,\lambda]}(u_{k})\Big)\,\textup{d}v\,\textup{d}y\,\textup{d}u\,\textup{d}x\,\frac{\textup{d}s}{s}\,\frac{\textup{d}t}{2\pi t},$
@@
$\times(\sigma_{a}\ast\mathbbm{g}_{r})(z-z^{\prime})\,\lambda^{-m+1}\Big(\prod_{k=1}^{m-1}\mathbbm{1}_{[-\lambda,-\theta\lambda]\cup[\theta\lambda,\lambda]}(u_{k})\Big)\,\textup{d}z\,\textup{d}z^{\prime}\,\textup{d}u\,\textup{d}x\,\frac{\textup{d}s}{s}\,\frac{\textup{d}t}{t}.$
@@
Due to our assumption on the exponential growth of the numbers $\lambda_{j}$, every positive number $s$ lies in a bounded number of intervals $\varepsilon\lambda_{j}/(10e),\lambda_{j}/10]$, $1\leqslant j\leqslant J$, with a bound depending on $\varepsilon$ alone. Therefore, applying the last estimate with $\lambda=\lambda_{j}$ and summing in $j$, we get
@@
$\times\Big(\prod_{\begin{subarray}{c}1\leqslant i\leqslant m\\ i\neq k\end{subarray}}\mathbbm{g}_{s\gamma_{i}}(x_{i}^{0}-x_{i}^{1})\Big)\,\textup{d}(x_{1}^{0},x_{1}^{1},x_{2}^{0},x_{2}^{1},\ldots,x_{m}^{0},x_{m}^{1})\,\frac{\textup{d}s}{s}$
for any $1\leqslant k\leqslant m$, any parameters $\gamma_{1},\ldots,\gamma_{m}\in(0,\infty)$, and a real-valued bounded and compactly supported measurable function $F$ on $\mathbb{R}^{d_{1}+d_{2}+\cdots+d_{m}}$. In our case, $d_{1}=\cdots=d_{m-1}=1$ and $d_{m}=n-m+1$, but the greater generality will only make the arguments more elegant and symmetric.
@@
$:=$ $\prod_{(r_{1},\ldots,r_{k-1},r_{k+1},\ldots,r_{n})\in\{0,1\}^{m-1}}F(x_{1}^{r_{1}},\ldots,x_{k-1}^{r_{k-1}},x,x_{k+1}^{r_{k+1}},\ldots,x_{m}^{r_{m}})$
@@
$0=\mathcal{N}_{\lambda_{j}}^{0}(A;R)\geqslant\frac{1}{3}\mathcal{N}_{\lambda_{j}}^{1}(A;R)>0$
@@ 10. Hyperbolic embeddings: proof of Theorem 10
Rotating the plane by $45^{\circ}$ and stretching it by factor $\sqrt{2}$ we can, more conveniently, replace the hyperbola $\mathscr{H}$ with the rotated hyperbola
@@
Here $u$ is a shorter notation for the $n$-tuple $(u_{1},\ldots,u_{n})$. As expected, $\mathcal{N}_{\lambda}^{0}(A;R)$ can again be obtained in the limit of $\mathcal{N}_{\lambda}^{\varepsilon}(A;R)$ as $\varepsilon\to 0$, which is shown easily, as in Lemma 8.1. For every $a,\lambda>0$ the support of the anisotropically dilated measure $\mu_{a\lambda,a\lambda^{-1}}$ lies on the hyperbola $\mathscr{H}^{\prime}$ stretched by factor $a$. Thus, $\mathcal{N}_{\lambda}^{0}(A;R)>0$ for any value of $\lambda>0$ clearly means that there exists the desired configuration (1.5) inside the set $A$.
@@
$0,\lceil R/\lambda\rceil\lambda]\times[0,\lceil R\lambda\rceil\lambda^{-1}]\supseteq[0,R]^{2}$
@@
$\times(\mu\ast\mathbbm{k}_{t})_{a_{j}\lambda,a_{j}\lambda^{-1}}(u_{j})\Big(\prod_{\begin{subarray}{c}1\leqslant k\leqslant n\\ k\neq j\end{subarray}}(\mu\ast\mathbbm{g}_{t})_{a_{k}\lambda,a_{k}\lambda^{-1}}(u_{k})\Big)\,\textup{d}u\,\textup{d}z\,\frac{\textup{d}t}{2\pi t}.$
@@
$\times(\mu\ast\mathbbm{k}_{t})_{a_{j}\lambda,a_{j}\lambda^{-1}}(u_{j})\Big(\prod_{\begin{subarray}{c}1\leqslant k\leqslant n\\ k\neq j\end{subarray}}(\mu\ast\mathbbm{g}_{t})_{a_{k}\lambda,a_{k}\lambda^{-1}}(u_{k})\Big)\,\textup{d}u\,\textup{d}z\,\frac{\textup{d}t}{2\pi t}.$
@@
$\qquad\qquad\quad\times(\mu\ast\mathbbm{g}_{t/\sqrt{2}})_{a_{j}\lambda,a_{j}\lambda^{-1}}(w-w^{\prime})\,\textup{d}w\,\textup{d}w^{\prime},$
@@
$\times(\mu\ast\mathbbm{g}_{t/\sqrt{2}})_{a_{j}\lambda,a_{j}\lambda^{-1}}(w-w^{\prime})\Big(\prod_{\begin{subarray}{c}1\leqslant k\leqslant n\\ k\neq j\end{subarray}}(\mu\ast\mathbbm{g}_{t})_{a_{k}\lambda,a_{k}\lambda^{-1}}(u_{k})\Big)\,\textup{d}w\,\textup{d}w^{\prime}\,\textup{d}(u_{k})_{k\neq j}\,\frac{\textup{d}t}{t}$
@@
$\times(\mu\ast\mathbbm{g}_{t/\sqrt{2}})_{a_{j}\lambda,a_{j}\lambda^{-1}}(w-w^{\prime})\Big(\prod_{\begin{subarray}{c}1\leqslant k\leqslant n\\ k\neq j\end{subarray}}(\mu\ast\mathbbm{g}_{t})_{a_{k}\lambda,a_{k}\lambda^{-1}}(u_{k})\Big)\,\textup{d}w\,\textup{d}w^{\prime}\,\textup{d}(u_{k})_{k\neq j}\,\frac{\textup{d}t}{t}.$
@@ 11. Remarks for harmonic analysts
One might wonder about the analytical reason why a density result for area $1$ rectangles is not possible in any dimension, while we were able to prove such a result for area 1 right-angled triangles.
Namely, take Schwartz functions $\varphi$ and $\psi$, respectively defined on $\mathbb{R}^{d_{1}}$ and $\mathbb{R}^{d_{2}}$, that have compact frequency supports and suppose that at least one of these supports does not contain the origin. Also, take some $4$-tuple $(p_{1},p_{2},p_{3},p_{4})$ of Lebesgue exponents in $1,\infty]$ satisfying the Hölder scaling,
@@
$\times\varphi_{t}(x-x^{\prime})\psi_{1/t}(y-y^{\prime})\,\textup{d}y\,\textup{d}y^{\prime}\,\textup{d}x\,\textup{d}x^{\prime}\,\frac{\textup{d}t}{t}\Big|\leqslant C_{r,R}\prod_{k=1}^{4}\|f_{k}\|_{\textup{L}^{p_{k}}(\mathbb{R}^{d_{1}+d_{2}})}$
grows like a multiple of $\log(R/r)$ as $R/r\to\infty$. This is not any better than in the trivial estimate obtained from Hölder’s inequality applied for each fixed $t$. Rather than giving a direct argument for this observation, one could argue that any nontrivial cancellation, i.e., a bound of the form $o(\log(R/r))$ as $R/r\to\infty$, would translate to a positive result for rectangles of area $1$. We would simply adjust the proof of Theorem 5 to work for fixed area rectangles, making it similar to the proof of Theorem 1, which we know to be impossible due to Theorems 3 and 4.
@@
so we only need to make sure that $\varphi$ and $\psi$ are such that the last integral is nonzero. The fact that the above quadrilinear form does not satisfy any nontrivial estimates is in sharp contrast with results for the so-called twisted paraproduct, where the scales are aligned as $\varphi_{t},\psi_{t}$, rather than $\varphi_{t},\psi_{1/t}$. It was shown to be bounded, at least when the exponents $p_{k}$ are greater than $2$, by Durcik [18, 19] and, in a dyadic model, previously also by the author [42].
@@
now actually has non-trivial cancellation over the scales $t$. Its best constants $C^{\prime}_{r,R}$ grow at most like $O((\log(R/r))^{1/2})$, which is easily seen similarly as in the proof of Lemma 8.4: by separating one of the functions and applying the classical square function bound to it. However, one could still ask if the above trilinear form is also bounded as $r\to 0+$ and $R\to\infty$, or, equivalently, if the constants $C^{\prime}_{r,R}$ are even independent of the scale truncation parameters $r$ and $R$. If, on the other hand, the scales were aligned as $\varphi_{t},\psi_{t}$, this would again be the bound for the twisted paraproduct and it would follow from [42, Theorem 1].
@@
12. Tabular summary
This section offers a condensed landscape of most (but not all) density results discussed in the present paper. Our goal is to emphasize distinctions among different point configurations and different types of “largeness” conditions imposed on the containing set.
[…]
minimal dim. increased dim. minimal dim. increased dim.
segment all A ([55])
large A [^1] A trivial one-dimensional set avoiding all half-integer distances is the union of intervals [k1/5,k+1/5][k-1/5,k+1/5] centered at integers $k$. U ([34, 32, 7]) A [^2] In any number of dimensions we can simply take a union of balls of small radius $\varepsilon>0$ around the points $(k,0,\ldots,0)$, $k\in\mathbb{Z}$.
right triangle U ([40])
simplex U ([24])
box all A (Thm. 4)
[…]
hypercube graph all A (Thm. 8) ? A (Thm. 8) ? [^3] For the particular case of the $1$-skeleton of the $2$-cube with prescribed product of edge-lengths $a_{1}a_{2}$, an avoiding set of infinite measure in $\mathbb{R}^{n}$ can be constructed as a union of balls around integer points $(k_{1},\ldots,k_{n})\in\mathbb{Z}^{n}$ with radii $\varepsilon/(|k_{1}|+\cdots+|k_{n}|+1)$, where $\varepsilon>0$ is sufficiently small.
large U ([45]) U ([48]) ?
hyperbolic segment U ([5], Thm. 9) A (Rem. 11)
hyperbolic star U (Thm. 10) A (Rem. 11)
Table 1. Landscape of results.
In Table 1 we discuss configurations that are vertex-sets of a segment, a right triangle, a simplex, a rectangular box, or a parallelotope, each of a fixed length/area/volume. Additionally, we consider a hypercube graph with a prescribed product of edge-lengths, a hyperbolic unit segment, i.e., a pair of points the difference of which lies on the standard hyperbola, or a hyperbolic star made of several hyperbolic segments. We mark whether they can be found in every set of positive upper density, or already in every set of infinite measure. When we say that a configuration is unavoidable (U), this means that it is found in all sets that are large in the appropriate sense. Otherwise a configuration is avoidable (A). Sometimes we make a further distinction by trying to find configurations that span polytopes of either all prescribed volumes or merely only sufficiently large ones. Moreover, different results are sometimes available depending on whether the ambient dimension is the minimal one that can host the configuration, or an increased one. The reader should understand that an isolated question mark (?) simply means that the author is not aware of any relevant literature, including the possibility that further investigation reveals the entry as being easy. The open problems listed in the introduction have already attracted interest, having been raised on several previous occasions.